AMMONIA CLUSTER DISSOCIATION VIA NH STRETCH EXCITATION

AMANDA S. CASE, \underline{CORNELIA G. HEID} and F. FLEMING CRIM, The University of Wisconsin-Madison, Department of Chemistry, 1101 University Avenue, Madison, WI 53706.

Ammonia clusters were formed in a supersonic expansion and interrogated with IR light in the NH stretching region. The ammonia molecules resulting from vibrational predissociation were detected using (2+1) resonance enhanced multiphoton ionization (REMPI) through the B̃ state. By monitoring various rotational levels of ammonia, we have also obtained an IR spectrum for the multiple clusters formed in our molecular beam.